OpenStudy (anonymous):

how come teacher dont use pemdas on equation like 4(3x+7)-2(8x-3)

OpenStudy (anonymous):

how else would they do it PEMDAS parenthesis- you cant add variables to constants, so we ignore this exponents- there are none multiplication- this can be done, you would have to distribute the 4 and distribute the -2 division- also none exists addition- add like terms subtraction- subtract like terms

OpenStudy (anonymous):

for example on a question 1/2(3x-8/6)-1/3(x-6/4)-(3x+3/4) the teacher got -11/6x-11/12. wouldn't it be 3.5/3x-11/12

OpenStudy (compassionate):

4(3x+7)-2(8x-3 12x + 28 - 16x + 6 -3x + 34

OpenStudy (anonymous):

er one sec 1/2(3x-8/6)-1/3(x-6/4)-(3x+3/4) 3/2x - 8/12 -1/3x +6/12 -3x -3/4 3/2x-2/3 -1/3x +1/2 -3x -3/4 9/6x -2/6x-18/6 x -8/12 +6/12 - 9/12 11/6 x -11/12

OpenStudy (anonymous):

oops sorry its -11/6x -11/12

OpenStudy (anonymous):

but aren't suppose to follow PEMDAS

OpenStudy (anonymous):

i did follow pemdas

OpenStudy (anonymous):

1/2(3x-8/6)-1/3(x-6/4)-(3x+3/4) starting with this P-i cant add the terms within the parenthesis so i leave the problem as is E- no exponents M- multiplication = distributing all the terms which would give me 3/2x - 8/12 -1/3x +6/12 -3x -3/4 D- no division then i just changed the fractions around so they have a common denominator which allows me to combine like terms easier 9/6x -2/6x-18/6 x -8/12 +6/12 - 9/12 then i did A & S - combined like terms -11/6x -11/12

OpenStudy (anonymous):

so, do you get it?

OpenStudy (anonymous):

this is how the teacher did it 1st she simplify to get 3/2x-8/12-1/3x=6/12-3x-3/4 then she simplify -1/3x and -3x together 1st since she told us it would be easier wouldn't it be 3/2x +-1/3 first

OpenStudy (anonymous):

oh, so your question has to do with order in which you add things it doesnt matter in what order you add things up its called the commutative property of addition and associative property of addition so the commutative property of addition lets us rearrange terms when you're adding/ subtracting them so A+B= B+A or A+B+C+D+E= A+B+C+E+D = A+B+D+E+C= E+D+C+B+A= ... etc now the associative property of addition says that it doesnt matter in which order you add multiple terms Associative property of addition (A+B)+C= A+(B+C) = (A+C)+B or

OpenStudy (anonymous):

ok thank you so much. I got so confuse

OpenStudy (anonymous):

yup no prob hopefully you understand this now